OpenStudy (anonymous):

If a gas has a volume of 1 L at a pressure of 270 kPa, what volume will it have when the pressure is increased to 540 kPa? Assume the temperature and number of particles are constant. Show your work

OpenStudy (stmoney):

Use the equation P1*V1 = P2*V2; T = temperature and P = Pressure. (270kPa)(1L) = (540 kPa)(V2). Solve for V2

OpenStudy (anonymous):

This is basically the same thing as what has already been posted...with a little bit more detail